A 25 Ω load is connected to the terminals of a 400 V terminal voltage parallel excited DC generator. The resistance of the field circuit is 200 Ω, the resistance of the armature is 0,4 Ω, the efficiency is 82%, the speed is 1000 r/min. Draw the equivalent circuit of the generator. Calculate the induced emf, the torque and the magnetic losses if the mechanical losses is 200 W.
Answer #1

252 Load, terminal voltage (4) - 4000 Resistance of field circuit IRA) = 2002 Resistance of armature circuit (Ra)= 0142 m%= 8Lossy=1404 - 87 watt = 18 2x x 2 + 2 ФЕРТПат сhic 2) P magnetic = 245024 watt so magnetic losses = 257.24 watt